OpenStudy (anonymous):

Starting with the definition 1.00in. = 2.54 cm, find the number of kilometers in 7.00mi . Please describe conversions.

OpenStudy (doc.brown):

You are starting out with two things that are equal. 1.00in is equal in length to 2.54cm. If you divide something by something that is equal you get 1. For example 2/2 = 1 or in this case 1.00in / 2.45cm = 1. A conversion doesn't change the size or length or mass or whatever, it only changes the units. For example 7in * 1 = 7in. or 7in * (2.54cm / 1.00in) = 17.78cm. If I were to tell you that 1 mile = 80 chains, 4 rods = 1 chain, 2 rods = 33 feet, 1 foot = 12 inches, 1 inch = 2.54 centimetres, 100 cm = 1 metre, 1 kilometre = 1000 metres, then: 7mi * (80chain / rod) * (4rod / chain) * (33ft / 2rod) * (12in / ft) * (2.54cm / in) * (m / 100cm) * (km / 1000m) = ( 7 * 80 * 4 * 33 * 12 * 2.54 * km) / (2 * 100 * 1000) = 11.3km